What is the voltage across a 3-μF capacitor if the charge on one plate is
0.12 mC? How much energy is stored

Grade:11

1 Answers

raju
8 years ago
(a) Since q = Cv,
q = 3 × 10 −12 × 20 = 60 pC
(b) The energy stored is
11w = Cv 2 = × 3 × 10 −12 × 400 = 600 pJ